OpenStudy (anonymous):

Suppose that y is directly proprtiona to x and that y=-55 when x=7. complete part B which is use y=kx o find when x=6.

OpenStudy (anonymous):

y=kx just put the 1st pair given to get k

OpenStudy (anonymous):

(-55,7)

OpenStudy (anonymous):

nothat is the answer for part A I need part B.

OpenStudy (anonymous):

did you find k

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the first part is to get k part A

OpenStudy (anonymous):

I got that In eedPart B

OpenStudy (anonymous):

-55=k*7

OpenStudy (anonymous):

It rought it wouldbe -55=y*6

OpenStudy (anonymous):

so k=-55/7 so we need y when x=3 so y=k*3 but we know k so??

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

not 3

OpenStudy (anonymous):

ifyou e cnfused read the end of the question

OpenStudy (anonymous):

yes you corrected that

OpenStudy (anonymous):

so what is y

OpenStudy (anonymous):

y=-55/7*6

OpenStudy (anonymous):

they said it was -330/7

OpenStudy (anonymous):

thanks fo your help